# 3D Dynamics-Aware Manipulation: Endowing Manipulation Policies with 3D Foresight

### 05 Mar 2026 v3

@@ I INTRODUCTION
An exciting direction for improving language-conditioned manipulation policies is the incorporation of world modeling, i.e., predicting the transition of world states driven by impetus like language command or low-level actions. Recently, this line of work has demonstrated promising results by pretraining policy models on large-scale video data to predict future RGB observation [30, 16, 25, 13, 34, 28]. Through this kind of 2D world model learning, policy models become aware of the desired future states under current observations and can more easily predict actions that lead to such states.
@@ II-A Language-conditioned Manipulation
Instead of constructing a single-task visuomotor policy for every task, it is more desirable to have a multitask language-conditioned manipulation policy that is generalizable. CLIPort [24], which injects frozen language embeddings into two-stream convolution networks, is one of the first studies that show the potential of this paradigm. However, purely relying on frozen language embeddings will impose a limit of generalizability, due to the domain gap between language and vision. Many methods [17, 5, 15, 3] leverage large vision language models (VLMs) to address this. Despite the promising results that these methods have achieved, they face efficiency and interpretability issues. Another line of work [30, 16, 25, 13, 34, 28] tries to mine the dynamics hidden in captioned videos, which is known as the world modeling approach [11].
@@ II-B World Modeling for Manipulation Policy Learning
Similar to VLM-based methods, the world modeling approach benefits from self-supervised learning on easily accessible unlabeled data. But the world modeling approach is more interpretable, as the core of it is to predict the outcome of behaviors. Existing work [30, 16, 25, 13, 34, 28] mainly leverages future image generation as the self-supervised learning objective. However, the dynamics of pixels in the 2D frame space is superficial and just loosely related to the dynamics of robotic actions. Our work highlights the significance of modeling the dynamics in 3D space.
@@ III METHOD
A language-conditioned action-chunking policy $\pi_{\theta}(\mathbf{a}_{t:t+K-1}|\mathbf{o}_{t-T+1:t},\mathbf{c})$ takes a language command $\mathbf{c}$ and observations $\mathbf{o}_{t-T+1:t}$ (images, proprioception states, etc.) with historical window size $T$ as input, and outputs a chunk of actions $\mathbf{a}_{t:t+K-1}$ of length $K$. These variables belong to quite distinct modalities, which is why knowledge from large-scale diverse data and auxiliary learning objectives that narrow the domain gap are necessary. ManiTrend incorporates the prediction of current depth $\mathbf{d}^{\text{main/wrist}}_{t}$, future RGB-D $\mathbf{rgbd}^{\text{main/wrist}}_{t+S}$ in main/wrist views and 3D flow $\boldsymbol{\tau}_{t:t+L-1}\in\mathbb{R}^{L\times P\times 3}$, where $S$ is the time shift of future images, $L$ is the flow length and $P$ is the number of track points. In our formulation, the three dimensions of 3D flow correspond to x, y (in pixel coordinates) and metric depth value, respectively. Our framework employs a causal transformer to model the multimodal spatial-temporal correlation, as shown in Fig. 2.
@@ V CONCLUSION
Taken together, we explore a way towards 3D dynamics-aware robotic manipulation in this paper. To enhance manipulation policies with 3D foresight, a novel framework is introduced, which integrates 3D world modeling and policy learning through three self-supervised objectives — current depth estimation, future RGB-D prediction, and 3D flow prediction. Experiments across simulation and the real world demonstrate that manipulation policies can benefit much more from 3D foresight than from 2D foresight. Our ablation study reveals the complementarity of the proposed self-supervised learning objectives and the major contribution of 3D dynamics-related objectives to performance. Our case study suggests that 3D foresight can provide manipulation policies with stronger spatial awareness which are critical for tasks that require distance perception or involve prominent depth-wise movement. A promising direction for future work is to explore more advanced 3D scene representations, such as point clouds, Tri-Plane or 3D Gaussian Splatting, to further enhance the model’s spatial reasoning capabilities.

### 26 Mar 2026 v4

@@ I INTRODUCTION
An exciting direction for improving language-conditioned manipulation policies is the incorporation of world modeling, i.e., predicting the transition of world states driven by impetus like language command or low-level actions. Recently, this line of work has demonstrated promising results by pretraining policy models on large-scale video data to predict future RGB observation [30, 16, 25, 34, 28]. Through this kind of 2D world model learning, policy models become aware of the desired future states under current observations and can more easily predict actions that lead to such states.
@@ II-A Language-conditioned Manipulation
Instead of constructing a single-task visuomotor policy for every task, it is more desirable to have a multitask language-conditioned manipulation policy that is generalizable. CLIPort [24], which injects frozen language embeddings into two-stream convolution networks, is one of the first studies that show the potential of this paradigm. However, purely relying on frozen language embeddings will impose a limit of generalizability, due to the domain gap between language and vision. Many methods [17, 5, 15, 3, 13] leverage large vision language models (VLMs) to address this. Despite the promising results that these methods have achieved, they face efficiency and interpretability issues. Another line of work [30, 16, 25, 34, 28] tries to mine the dynamics hidden in captioned videos, which is known as the world modeling approach [11].
@@ II-B World Modeling for Manipulation Policy Learning
Similar to VLM-based methods, the world modeling approach benefits from self-supervised learning on easily accessible unlabeled data. But the world modeling approach is more interpretable, as the core of it is to predict the outcome of behaviors. Existing work [30, 16, 25, 34, 28] mainly leverages future image generation as the self-supervised learning objective. However, the dynamics of pixels in the 2D frame space is superficial and just loosely related to the dynamics of robotic actions. Our work highlights the significance of modeling the dynamics in 3D space.
@@ III METHOD
A language-conditioned action-chunking policy $\pi_{\theta}(\mathbf{a}_{t:t+K-1}|\mathbf{o}_{t-T+1:t},\mathbf{c})$ takes a language command $\mathbf{c}$ and observations $\mathbf{o}_{t-T+1:t}$ (images, proprioception states, etc.) with historical window size $T$ as input, and outputs a chunk of actions $\mathbf{a}_{t:t+K-1}$ of length $K$. These variables belong to quite distinct modalities, which is why knowledge from large-scale diverse data and auxiliary learning objectives that narrow the domain gap are necessary. 3D Foresight incorporates the prediction of current depth $\mathbf{d}^{\text{main/wrist}}_{t}$, future RGB-D $\mathbf{rgbd}^{\text{main/wrist}}_{t+S}$ in main/wrist views and 3D flow $\boldsymbol{\tau}_{t:t+L-1}\in\mathbb{R}^{L\times P\times 3}$, where $S$ is the time shift of future images, $L$ is the flow length and $P$ is the number of track points. In our formulation, the three dimensions of 3D flow correspond to x, y (in pixel coordinates) and metric depth value, respectively. Our framework employs a causal transformer to model the multimodal spatial-temporal correlation, as shown in Fig. 2.
@@ V CONCLUSION
We explore a way towards 3D dynamics-aware robotic manipulation in this paper. To enhance manipulation policies with 3D foresight, a novel framework is introduced, which integrates 3D world modeling and policy learning through three auxiliary objectives — current depth estimation, future RGB-D prediction, and 3D flow prediction. Experiments across simulation and the real world demonstrate that policies can benefit much more from 3D foresight than from 2D foresight. Our ablation study reveals the complementarity of the proposed learning objectives and the major contribution of 3D dynamics-related objectives to performance. Our case study suggests that 3D foresight can provide manipulation policies with stronger spatial awareness which are critical for tasks that require distance perception or involve prominent depth-wise movement. A promising direction for future work is to explore more advanced 3D scene representations to further enhance the model’s spatial reasoning capabilities.
VI ACKNOWLEDGMENT
This work is in part supported by the Guangzhou-HKUST(GZ) Joint Funding Program (2025A03J3656) , in part supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2026A1515012291).